OpenStudy (anonymous):

Given a = -1, h = 3, and k = 2, how will the graph g(x) change if the parent function is f(x) = x2?

OpenStudy (lynfran):

g(x)=a(x+h)+k g(x)=-1(x-3)+2 since a is negative this graph opens downwards while the parent function look like this it opens upward

OpenStudy (anonymous):

The graph will be reflected and move 2 units down. The graph will move 3 units to the left and 2 units up. The graph will move 3 units to the right and 2 units up. The graph will be reflected, it will move 3 units to the right, and 2 units up.